OpenStudy (jfruscianterhcp):

The diagonal of a square is 12 units. find the are of the square

OpenStudy (jfruscianterhcp):

area*

OpenStudy (anonymous):

if the diagonal is \(a\) then the side is \(\frac{a}{\sqrt{2}}\) that makes the area \[\left(\frac{a}{\sqrt{2}}\right)^2=\frac{a^2}{2}\]